OpenStudy (anonymous):

Z^15 * Z^0 = Z^15 correct

OpenStudy (turingtest):

yes

Directrix (directrix):

Corrrect. Z^0 = 1 and that leaves Z^15 *1 = Z^15

OpenStudy (turingtest):

...unless z=0

OpenStudy (anonymous):

no you were right

OpenStudy (turingtest):

I'm just pointing out that this is only true if we assume that \(z\neq0\) because \(0^0\) is undefined.